OpenStudy (anonymous):

what is the square root of 4/49 show work

OpenStudy (anonymous):

You can rewrite the expression \[\sqrt{(\frac{4}{49})}\] Like this: \[\frac{\sqrt{4}}{\sqrt{49}}\]

OpenStudy (anonymous):

then evaluate that...

OpenStudy (anonymous):